OpenStudy (anonymous):

SOLVE FOR m m/30 = 10/50

OpenStudy (abdullahm):

cross multiply

OpenStudy (abdullahm):

\(\sf\Large\frac{a}{b} = \frac{c}{d} \to ad =bc\)

OpenStudy (anonymous):

so you get m= 300/50

OpenStudy (anonymous):

which is m=6